# Adapt and Diffuse: Sample-adaptive Reconstruction via Latent Diffusion Models

### v2

@@ Appendix A Experimental details
DPS [5]: DPS approximates the gradient of the likelihood as
@@
$\bm{z}_{i-1}=\bm{z^{\prime}}_{i-1}-\eta_{i}\nabla_{\bm{z}_{i}}\left\|\mathcal{A}(\hat{\bm{z}}_{0}(\bm{z}_{i})-\bm{y}\right\|^{2}.$ (7)
@@
GML-DPS [34]: Latent-DPS achieves poor performance out-of-the-box due to the inaccuracy of the approximation in 6 in early stages of diffusion and due to additional complications introduced by the decoder [34, 39, 9]. GML-DPS is an extension of Latent-DPS that guides the diffusion towards fixed points of the encoder-decoder composition via the "goodness modified" approximation
$\nabla_{\bm{z}_{i}}\log p(\bm{y}|\bm{z}_{i})\approx\nabla_{\bm{z}_{i}}\log p(\bm{y}|\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i})))+\lambda\nabla_{\bm{z}_{i}}\|\hat{\bm{z}}_{0}(\bm{z}_{i})-\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i}))\|^{2},$ (9)
@@
$\bm{z}_{i-1}=\bm{z^{\prime}}_{i-1}-\eta_{i}^{LDPS}\nabla_{\bm{z}_{i}}\left\|\mathcal{A}(\hat{\bm{z}}_{0}(\bm{z}_{i})-\bm{y}\right\|^{2}-\gamma_{i}^{GM}\nabla_{\bm{z}_{i}}\|\hat{\bm{z}}_{0}(\bm{z}_{i})-\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i}))\|^{2},$ (10)
@@
$\gamma_{i}^{GM}=\frac{\gamma^{GM}}{\left\|\hat{\bm{z}}_{0}(\bm{z}_{i})-\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i}))\right\|},$ (11)
@@
PSLD [34]: PSLD improves upon GML-DPS by adding a "gluing" term in order to avoid inconsistencies at mask boundaries. The approximation takes the form
$\nabla_{\bm{z}_{i}}\log p(\bm{y}|\bm{z}_{i})\approx\nabla_{\bm{z}_{i}}\log p(\bm{y}|\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i})))+\lambda\nabla_{\bm{z}_{i}}\|\hat{\bm{z}}_{0}(\bm{z}_{i})-\mathcal{D}_{0}(\mathcal{A}\bm{x}_{0}^{*}+(\mathbf{I}-\mathcal{A}\mathcal{A}^{T})\hat{\bm{z}}_{0}(\bm{z}_{i}))\|^{2},$ (12)
@@
$\bm{z}_{i-1}=\bm{z^{\prime}}_{i-1}-\eta_{i}^{LDPS}\nabla_{\bm{z}_{i}}\left\|\mathcal{A}(\hat{\bm{z}}_{0}(\bm{z}_{i})-\bm{y}\right\|^{2}-\gamma_{i}^{PSLD}\nabla_{\bm{z}_{i}}\|\hat{\bm{z}}_{0}(\bm{z}_{i})-\mathcal{D}_{0}(\mathcal{A}\bm{x}^{*}+(\mathbf{I}-\mathcal{A}\mathcal{A}^{T})\hat{\bm{z}}_{0}(\bm{z}_{i}))\|^{2},$ (13)
@@
$\gamma_{i}^{PSLD}=\frac{\gamma^{PSLD}}{\left\|\hat{\bm{z}}_{0}(\bm{z}_{i})-\mathcal{D}_{0}(\mathcal{A}\bm{x}^{*}+(\mathbf{I}-\mathcal{A}\mathcal{A}^{T})\hat{\bm{z}}_{0}(\bm{z}_{i}))\right\|},$ (14)
and $\gamma^{PSLD}$ is a hyperparameter. Note that PSLD requires access to noiseless measurements $\mathcal{A}\bm{x}^{*}$ and to the transposed of the linear operator $\mathcal{A}$. In our experiments that involve measurement noise, we replace $\mathcal{A}\bm{x}^{*}$ with $\bm{y}$ (noisy measurement). For Gaussian blur, we implement the transposed operator $\mathcal{A}^{T}$ using ConvTranspose2d in torch. We omit experiments involving nonlinear blur with PSLD as the transposed operator is not known.
@@
ReSample [39]: Instead of DPS-like approximations as in the previous solvers, ReSample applies data consistency directly on the posterior mean estimates by solving the latent-domain optimization problem
@@
or its pixel-domain variant after passing $\hat{\bm{z}}_{0}(\bm{z}_{i})$ through the decoder. To solve the above problem, numerical methods are used where optimization is terminated when the stopping condition $\frac{1}{m}\|\bm{y}-\mathcal{A}(\mathcal{D}_{0}(\bm{z}))\|_{2}^{2}<\tau$ is reached. $\tau$ is a hyperparameter, which we set to $\sigma_{y}^{2}$ as reducing the error below the noise level amounts to overfitting to the noisy measurement. Furthermore, we set $\lambda=1$ for all experiments following [39], as we observe that the method is not sensitive to the choice of this parameter as long as the problem is solved to the desired accuracy. We use Adam optimizer with step size $0.1$ and initial guess $\hat{\bm{z}}_{0}(\bm{z}_{i})$ to minimize the objective in (15). As solving (15) is computationally heavy, it is only performed every $K$ steps, where we set $K=10$ for all experiments following the setting in the original paper.
@@
where only $\omega$ needs to be tuned. We do not use time-travel updates as in the original algorithm as it appears to have only minor contribution to performance (see Table 8. in [39]). Furthermore, we use DDIM sampling with $N=500$ matching the setting in the original paper.
@@ Appendix B Ablation of Flash-Diffusion components
Appendix C Experiment on contributors to degradation severity
Here, we analyze the contribution of different factors to the predicted degradation severities (Fig. 8). We apply severity encoding to both the clean image with noise (no blur) and the noisy and degraded image, resulting in predicted severities $\hat{\sigma}_{noisy}$ and $\hat{\sigma}_{degr.+noisy}$. We quantify the difficulty of samples relative to each other via percentiles of the above two quantities, where we use $\hat{\sigma}_{noisy}$ as a proxy for the difficulty originating from the image structure. We observe that for a fixed $\hat{\sigma}_{degr.+noisy}$, the composition of degradation severity may greatly vary. The two presented images have been assigned approximately the same $\hat{\sigma}_{degr.+noisy}$, however the top image is inherently easy to encode (low $\hat{\sigma}_{noisy}$ percentile) compared to other images in the dataset, therefore the severity is mostly attributed to the image degradation. On the other hand, the bottom image with the same $\hat{\sigma}_{degr.+noisy}$ is less corrupted by blur, but with high $\hat{\sigma}_{noisy}$ indicating a difficult image. This example further corroborates the interaction between ground truth signal structure and the applied corruption in determining the difficulty of a reconstruction task.
Figure 8: Contributions to predicted severity. Degraded images with approx. the same $\hat{\sigma}$ may have different factors contributing to the predicted severity. The main contributor to $\hat{\sigma}$ in the top image is the image degradation (blur), whereas the bottom image is inherently difficult to reconstruct.
Appendix D Robustness studies
D.1 Noise robustness
@@
Our most crucial expectation towards the severity encoder is to predict a degradation severity $\hat{\sigma}(\bm{y})$ that is a non-decreasing function of the true underlying corruption level. More specifically, assume that the forward model $\mathcal{A}$ is parameterized by a severity level $s\in[0,1]$, with $s=0$ being least severe, $s=1$ being most severe. For instance, the standard deviation of the Gaussian blur kernel may serve as $s$ in case of Gaussian blur (see a more rigorous treatment of degradation severity in [12]). Then for a given specific clean image $\bm{x}_{0}$ and its corrupted versions $\bm{y}_{s^{\prime}}:=\mathcal{A}(\bm{x}_{0};s^{\prime})+\bm{n},~{}~{}\bm{y}_{s^{\prime\prime}}:=\mathcal{A}(\bm{x}_{0};s^{\prime\prime})+\bm{n}$, we expect our severity encoder to satisfy
@@
[…]
@@
Robustness of Flash-Diffusion performance –Next, we investigate the impact of noise level mismatch in test-time on the reconstruction performance of adaptive solvers imbued with Flash-Diffusion. In particular, we perform experiments on the FFHQ dataset with varying Gaussian blur and fixed measurement noise of $\sigma_{\bm{y}}=0.05$ in severity encoder training. We pick ReSample as our baseline solver as it yields the best performance among non-adaptive solvers in our experiments. As the stopping criteria for hard data consistency ($\tau$ in [39]) depends on the measurement noise, we expect degraded performance when the true noise level differs from the one used to tune the hyperparameters. We vary the measurement noise in test time and compare the reconstruction performance of Flash(ReSample) to ReSample and SwinIR. We depict the results of the robustness study in Figure 7. We observe that both in terms of perceptual and distortion metrics, the performance of Flash(ReSample) degrades more gracefully compared to its non-adaptive counterpart and SwinIR in most cases (with the exception of lower than expected noise and distortion metrics). We hypothesize that this additional robustness is due to the sample-by-sample adaptivity of Flash-Diffusion. As the noise level increases, our method can automatically adapt by increasing the number of diffusion steps due to higher degradation severity predictions from the severity encoder.
@@
D.2 Robustness against forward model mismatch
@@
Robustness of severity encoding – First, following our methodology in Appendix D.1, we investigate how severity encoding performance degrades if the image degradation forward model is different from what the severity encoder has been trained on. We train the severity encoder on nonlinear blur and evaluate it on Gaussian blur of varying amounts on the CelebA dataset. We expect the predicted severity to be a non-decreasing function of the true blur amount. However, as Figure 12 demonstrates, the monotonicity requirement is violated due to the forward model mismatch. In particular, the severity encoder fails to predict the degradation severities of lightly blurred images. However, we observe satisfactory performance at high blur amounts (above $0.3$ in Fig. 12). We hypothesize that at high blur levels, Gaussian blur and nonlinear blur become similar, thus the severity encoder trained on either may provide acceptable predictions for both. We further support this hypothesis by evaluating the ordering accuracy on a validation set of $100$ CelebA images. The results are summarized in Table 4. We observe that the ordering accuracy is poor (mean $\text{OA}=0.7704$) when evaluated across the full range of blur levels. However, it is near-perfect when low blur levels are excluded from the evaluation (mean $\text{OA}=0.9778$). This experiment highlights the severity encoder’s potential for generalization to unseen forward models, which opens up interesting future directions for fast adaptation to new tasks.
@@
Appendix E Experiments with accelerated baseline solvers
DDIM [40], an accelerated sampling scheme, has demonstrated comparable generation performance to DDPM samplers with significantly lower compute requirements, often able to reduce the number of reverse diffusion steps from $N=1000$ to $N=100$ in the context of image synthesis. However, to the best of our knowledge, similar acceleration has not been successfully translated to latent domain diffusion solvers without significant drop in reconstruction performance. In this section, we investigate:
@@ Appendix E Experiments with accelerated baseline solvers
$\bm{z}_{i-1}=\sqrt{\bar{\alpha}_{i-1}}\hat{\bm{z}}_{0}(\bm{z}_{t})+\sqrt{1-\bar{\alpha}_{i-1}-\delta^{2}\tilde{\beta}_{i}^{2}}\epsilon_{\bm{\theta}}(\bm{z}_{i},i)+\delta\tilde{\beta}_{i}^{2}\bm{\epsilon},$

### v3

@@ Appendix A Experimental Details
DPS (Chung et al., 2022a): DPS approximates the gradient of the likelihood as
@@
$\bm{z}_{i-1}=\bm{z^{\prime}}_{i-1}-\eta_{i}\nabla_{\bm{z}_{i}}\left\|\mathcal{A}(\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i}))-\bm{y}\right\|^{2}.$ (7)
@@
GML-DPS (Rout et al., 2024): Latent-DPS achieves poor performance out-of-the-box due to the inaccuracy of the approximation in 6 in early stages of diffusion and due to additional complications introduced by the decoder Rout et al. (2024); Song et al. (2023); Chung et al. (2023). GML-DPS is an extension of Latent-DPS that guides the diffusion towards fixed points of the encoder-decoder composition via the "goodness modified" approximation
$\nabla_{\bm{z}_{i}}\log p(\bm{y}|\bm{z}_{i})\approx\nabla_{\bm{z}_{i}}\log p(\bm{y}|\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i})))+\lambda\nabla_{\bm{z}_{i}}\|\hat{\bm{z}}_{0}(\bm{z}_{i})-\mathcal{E}_{0}(\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i})))\|^{2},$ (9)
@@
$\bm{z}_{i-1}=\bm{z^{\prime}}_{i-1}-\eta_{i}^{LDPS}\nabla_{\bm{z}_{i}}\left\|\mathcal{A}(\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i})))-\bm{y}\right\|^{2}-\gamma_{i}^{GM}\nabla_{\bm{z}_{i}}\|\hat{\bm{z}}_{0}(\bm{z}_{i})-\mathcal{E}_{0}(\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i})))\|^{2},$ (10)
@@
$\gamma_{i}^{GM}=\frac{\gamma^{GM}}{\left\|\hat{\bm{z}}_{0}(\bm{z}_{i})-\mathcal{E}_{0}(\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i})))\right\|},$ (11)
@@
PSLD (Rout et al., 2024): PSLD improves upon GML-DPS by adding a "gluing" term in order to avoid inconsistencies at mask boundaries. The approximation takes the form
$\nabla_{\bm{z}_{i}}\log p(\bm{y}|\bm{z}_{i})\approx\nabla_{\bm{z}_{i}}\log p(\bm{y}|\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i})))+\lambda\nabla_{\bm{z}_{i}}\|\hat{\bm{z}}_{0}(\bm{z}_{i})-\mathcal{E}_{0}(\mathcal{A}\bm{x}_{0}+(\mathbf{I}-\mathcal{A}\mathcal{A}^{T})\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i})))\|^{2},$ (12)
@@
$\bm{z}_{i-1}=\bm{z^{\prime}}_{i-1}-\eta_{i}^{LDPS}\nabla_{\bm{z}_{i}}\left\|\mathcal{A}(\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i})))-\bm{y}\right\|^{2}-\gamma_{i}^{PSLD}\nabla_{\bm{z}_{i}}\|\hat{\bm{z}}_{0}(\bm{z}_{i})-\mathcal{E}_{0}(\mathcal{A}\bm{x}_{0}+(\mathbf{I}-\mathcal{A}\mathcal{A}^{T})\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i})))\|^{2},$ (13)
@@
$\gamma_{i}^{PSLD}=\frac{\gamma^{PSLD}}{\left\|\hat{\bm{z}}_{0}(\bm{z}_{i})-\mathcal{E}_{0}(\mathcal{A}\bm{x}_{0}+(\mathbf{I}-\mathcal{A}\mathcal{A}^{T})\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i})))\right\|},$ (14)
and $\gamma^{PSLD}$ is a hyperparameter. Note that PSLD requires access to noiseless measurements $\mathcal{A}\bm{x}_{0}$ and to the transposed of the linear operator $\mathcal{A}$. In our experiments that involve measurement noise, we replace $\mathcal{A}\bm{x}_{0}$ with $\bm{y}$ (noisy measurement). For Gaussian blur, we implement the transposed operator $\mathcal{A}^{T}$ using ConvTranspose2d in torch. We omit experiments involving nonlinear blur with PSLD as the transposed operator is not known.
@@
ReSample (Song et al., 2023): Instead of DPS-like approximations as in the previous solvers, ReSample applies data consistency directly on the posterior mean estimates by solving the latent-domain optimization problem
@@
or its pixel-domain variant after passing $\hat{\bm{z}}_{0}(\bm{z}_{i})$ through the decoder. To solve the above problem, numerical methods are used where optimization is terminated when the stopping condition $\frac{1}{m}\|\bm{y}-\mathcal{A}(\mathcal{D}_{0}(\bm{z}))\|_{2}^{2}<\tau$ is reached. $\tau$ is a hyperparameter, which we set to $\sigma_{y}^{2}$ as reducing the error below the noise level amounts to overfitting to the noisy measurement. Furthermore, we set $\lambda=1$ for all experiments following Song et al. (2023), as we observe that the method is not sensitive to the choice of this parameter as long as the problem is solved to the desired accuracy. We use Adam optimizer with step size $0.1$ and initial guess $\hat{\bm{z}}_{0}(\bm{z}_{i})$ to minimize the objective in (15). As solving (15) is computationally heavy, it is only performed every $K$ steps, where we set $K=10$ for all experiments following the setting in the original paper.
@@
where only $\omega$ needs to be tuned. We do not use time-travel updates as in the original algorithm as it appears to have only minor contribution to performance (see Table 8. in Song et al. (2023)). Furthermore, we use DDIM sampling with $N=500$ matching the setting in the original paper.
@@ Appendix B Ablation of Flash-Diffusion Components
Appendix C Robustness Studies
C.1 Noise robustness
@@
Our most crucial expectation towards the severity encoder is to predict a degradation severity $\hat{\sigma}(\bm{y})$ that is a non-decreasing function of the true underlying corruption level. More specifically, assume that the forward model $\mathcal{A}$ is parameterized by a severity level $s\in[0,1]$, with $s=0$ being least severe, $s=1$ being most severe. For instance, the standard deviation of the Gaussian blur kernel may serve as $s$ in case of Gaussian blur (see a more rigorous treatment of degradation severity in Fabian et al. (2023)). Then for a given specific clean image $\bm{x}_{0}$ and its corrupted versions $\bm{y}_{s^{\prime}}:=\mathcal{A}(\bm{x}_{0};s^{\prime})+\bm{n},~{}~{}\bm{y}_{s^{\prime\prime}}:=\mathcal{A}(\bm{x}_{0};s^{\prime\prime})+\bm{n}$, we expect our severity encoder to satisfy
@@
$g(\bm{y}_{s_{i}},\bm{y}_{s_{j}};\bm{x}_{0})=\begin{cases}1&\text{if}~{}~{}~{}\text{sign}(s_{i}-s_{j})=\text{sign}(\hat{\sigma}_{i}-\hat{\sigma}_{j})\\ 0&\text{otherwise}\end{cases}$ (19)
@@
Robustness of Flash-Diffusion performance –Next, we investigate the impact of noise level mismatch in test-time on the reconstruction performance of adaptive solvers imbued with Flash-Diffusion. In particular, we perform experiments on the FFHQ dataset with varying Gaussian blur and fixed measurement noise of $\sigma_{\bm{y}}=0.05$ in severity encoder training. We pick ReSample as our baseline solver as it yields the best performance among non-adaptive solvers in our experiments. As the stopping criteria for hard data consistency ($\tau$ in Song et al. (2023)) depends on the measurement noise, we expect degraded performance when the true noise level differs from the one used to tune the hyperparameters. We vary the measurement noise in test time and compare the reconstruction performance of Flash(ReSample) to ReSample and SwinIR. We depict the results of the robustness study in Figure 8. We observe that both in terms of perceptual and distortion metrics, the performance of Flash(ReSample) degrades more gracefully compared to its non-adaptive counterpart and SwinIR in most cases (with the exception of lower than expected noise and distortion metrics). We hypothesize that this additional robustness is due to the sample-by-sample adaptivity of Flash-Diffusion. As the noise level increases, our method can automatically adapt by increasing the number of diffusion steps due to higher degradation severity predictions from the severity encoder.
@@
C.2 Robustness against forward model mismatch
@@
Robustness of severity encoding – First, following our methodology in Appendix C.1, we investigate how severity encoding performance degrades if the image degradation forward model is different from what the severity encoder has been trained on. We train the severity encoder on nonlinear blur and evaluate it on Gaussian blur of varying amounts on the CelebA dataset. We expect the predicted severity to be a non-decreasing function of the true blur amount. However, as Figure 12 demonstrates, the monotonicity requirement is violated due to the forward model mismatch. In particular, the severity encoder fails to predict the degradation severities of lightly blurred images. However, we observe satisfactory performance at high blur amounts (above $0.3$ in Fig. 12). We hypothesize that at high blur levels, Gaussian blur and nonlinear blur become similar, thus the severity encoder trained on either may provide acceptable predictions for both. We further support this hypothesis by evaluating the ordering accuracy on a validation set of $100$ CelebA images. The results are summarized in Table 4. We observe that the ordering accuracy is poor (mean $\text{OA}=0.7704$) when evaluated across the full range of blur levels. However, it is near-perfect when low blur levels are excluded from the evaluation (mean $\text{OA}=0.9778$). This experiment highlights the severity encoder’s potential for generalization to unseen forward models, which opens up interesting future directions for fast adaptation to new tasks.
@@
Appendix D Experiments with Accelerated Baseline Solvers
DDIM (Song et al., 2021a), an accelerated sampling scheme, has demonstrated comparable generation performance to DDPM samplers with significantly lower compute requirements, often able to reduce the number of reverse diffusion steps from $N=1000$ to $N=100$ in the context of image synthesis. However, to the best of our knowledge, similar acceleration has not been successfully translated to latent domain diffusion solvers without significant drop in reconstruction performance. In this section, we investigate:
@@ Appendix D Experiments with Accelerated Baseline Solvers
$\bm{z}_{i-1}=\sqrt{\bar{\alpha}_{i-1}}\hat{\bm{z}}_{0}(\bm{z}_{t})+\sqrt{1-\bar{\alpha}_{i-1}-\delta^{2}\tilde{\beta}_{i}^{2}}\epsilon_{\bm{\theta}}(\bm{z}_{i},i)+\delta\tilde{\beta}_{i}^{2}\bm{\epsilon},$ (21)
@@
Appendix E Adaptive DDIM Encoding Initialization
So far we have initialized the sampling trajectory with a noise-corrected latent estimate obtained from our severity encoder, where the noise correction is aimed at suppressing structure in the estimation error. This initialization scheme is analogous to stochastic encoding (Song et al., 2023). In this section, we investigate further techniques to initialize the reverse diffusion process in our framework in a sample-adaptive fashion.
With DDIM, the generative process can be reversed deterministically by simply running
$\bm{z}_{i+1}=\sqrt{\alpha_{i}}\bm{z}_{i}-\sqrt{1-\alpha_{i}}\sqrt{1-\bar{\alpha}_{i}}s_{\bm{\theta}}(\bm{z}_{i},i),$ (22)
which can be obtained from reversing (21). Here, $s_{\bm{\theta}}$ denotes the learned score (scaled version of $\epsilon_{\bm{\theta}}$). When starting from the clean latent $\bm{z}_{0}$ and running (22) for $t=0,..,N-1$ one obtains a deterministic encoding of $\bm{z}_{0}$ and the process is referred to as DDIM encoding (Preechakul et al., 2022). STSL (Rout et al., 2023) leverages DDIM encoding of a naive reconstruction ($\mathcal{E}_{0}(\bm{A}^{T}\bm{y})$, where $\varepsilon_{0}$ denotes the original encoder pretrained with the LDM) to obtain an initialization, where the DDIM forward process is run from $t=0$ to $T$. This initialization scheme can be combined with Flash adaptation, that is we can perform “partial” DDIM encoding from $t=0$ to $t_{start}$, where $t_{start}$ is the starting time index predicted by severity encoding. We refer to this technique as adaptive DDIM encoding. We perform experiments for the following scenarios:
  • STSL initialization with adaptation: we use adaptive DDIM encoding of $\mathcal{E}_{0}(\bm{A}^{T}\bm{y})$ by running the DDIM forward process for $t_{start}$ steps, where $t_{start}$ is obtained from the severity encoder. We do not use noise correction.
  • Flash initialization with DDIM encoding: we use adaptive DDIM encoding of $\hat{z}$ (latent estimate from severity encoder) by running the DDIM forward process for $t_{start}$ steps, where $t_{start}$ is obtained from the severity encoder. We do not use noise correction.
We choose Latent-DPS with DDIM sampling ($N=500$, $\eta=0.0$) as our baseline solver (same setting as described in Appendix D) in order to make the reverse process as close to DDIM sampling as possible within our framework. We show results on the FFHQ test set with varying Gaussian blur and random inpainting operators in Table 9. The mean $t_{start}$ predicted by severity encoding on the dataset is $49.91$ for deblurring and $52.08$ for inpainting.
Gaussian Deblurring (varying) Random Inpainting
Initialization SSIM()($\uparrow$)( ↑ ) PSNR()($\uparrow$)( ↑ ) LPIPS()($\downarrow$)( ↓ ) FID()($\downarrow$)( ↓ ) SSIM()($\uparrow$)( ↑ ) PSNR()($\uparrow$)( ↑ ) LPIPS()($\downarrow$)( ↓ ) FID()($\downarrow$)( ↓ )
$\hat{\bm{z}}$ + noise correction (ours) 0.8136 28.96 0.2328 56.91 0.8447 29.11 0.2010 57.24
$\hat{\bm{z}}$ + DDIM encoding 0.8488 30.05 0.2497 82.54 0.8605 29.38 0.2132 68.23
$\mathcal{E}_{0}(\bm{A}^{T}\bm{y})$ + DDIM encoding 0.7906 28.12 0.3201 92.08 0.6337 24.87 0.4302 135.4
Table 9: Results on the FFHQ dataset using various initialization schemes.
First, we observe that STSL initialization with adaptation performs worse than our vanilla Flash initialization. This is natural, as we leverage a supervised model to obtain an initial reconstruction that is of better quality than the unsupervised $\mathcal{E}_{0}(\bm{A}^{T}\bm{y})$ initialization. Furthermore, we suspect that since the autoencoder of LDM has been pretrained predominantly on clean images, it may compress $\bm{A}^{T}\bm{y}$ inefficiently.
Second, we find that adaptive DDIM encoding on top of the latent estimate $\hat{\bm{z}}$ obtained from our severity encoder performs surprisingly well. Even though this initialization falls short in terms of perceptual quality compared to vanilla Flash initialization with noise correction, it achieves superior distortion (PSNR, SSIM). This hints at blurrier, but highly reliable reconstructions faithful to the original measurement. This experiment suggests that Flash-Diffusion combined with adaptive DDIM encoding initialization may be very promising. In particular, by tuning a hyperparameter that controls the number of DDIM forward steps for encoding (similar to our noise correction parameter $c$) these results may be further improved. We leave this exploration for future work.
We note that a negative side-effect of using adaptive DDIM encoding in our setting is that it may limit the baseline solver to DDIM variants, as it is unclear how this initialization scheme performs when combined with arbitrary sampling techniques.